OpenStudy (anonymous):

find slope and y-intercept for 7x + 3y =50

OpenStudy (unklerhaukus):

when the equation is in this form \[y=mx+b\] \(m\) is the gradient and \(b\) is the \(y\)-intercept

OpenStudy (unklerhaukus):

to get your equation into this form , minus seven x form both sides and then divide both sides by three

OpenStudy (anonymous):

X=1/y=14

OpenStudy (unklerhaukus):

\[7x + 3y =50\] minus seven x \[7x + 3y -7x=50-7x\] \[3y=50-7x\] divide by 3 \[3y\div 3=(50-7x)\div 3\] \[\qquad\quad=\]